OpenStudy (anonymous):

how do i solve for x in the following, 3/7x-1/2=2/3

OpenStudy (anonymous):

3/7x-1/2=2/3 \[x = \frac{ 18 }{ 49 }\] http://www.themathpage.com/alg/equations-fractions.htm

OpenStudy (wwe123):

3/7x=2/3=1/2 3/7x=7/6 x=18/49

OpenStudy (anonymous):

Thank you, I needed to know how

OpenStudy (wwe123):

3/7x=2/3=1/2 3/7x=4+3/6 3/7x=7/6 7*7x=18 x=18/49

OpenStudy (anonymous):

There that helps. thank you

OpenStudy (wwe123):

so click on best response plz